how do I solve this word problem
"a rectangular rug has area of 96 square feet. If its leng is 4 feet more than twice its with. find the dimension of the rug. has to use algebratic equatio"
I first tried writing out the problem as this
L=4+2w but that is as far as I knew how to get.

Answer by rothauserc(4718)

we are given:
"a rectangular rug has area of 96 square feet. If its leng is 4 feet more than twice its with. find the dimension of the rug. has to use algebratic equatio"
you are nearly there :-)
l = 2w+4 and
(2w+4)*w=96 (area of a rectangle)
then we have
2w^2+4w-96=0
divide both sides of = sign by 2 and we get
w^2+2w-48=0
the above factors into
(w+8)*(w-6)=0
and we have two solutions for w, name -8 and 6
so w=6 and l=2*6+4=16